# Relaxation Fluctuations of Correlation Functions: Spin and Random Matrix Models

### v2

@@ I Introduction
With the motivation to study quantum chaotic systems and possibly to classify them, similar to their classical counterpart, we utilize the ideas put forth in Lakshminarayan and Balazs (1994); Lakshminarayan (1997, 1998) regarding the relaxation fluctuations of the correlation function and its intrinsic connection to chaos. The study was motivated by and studied using models having well-defined classical limits which are known to be chaotic such as in quantum maps Haake (1991); Lakshminarayan (2018); Balazs and Voros (1989); Lakshminarayan and Balazs (1994); Lakshminarayan (1997, 1998). By studying the quantum correlation functions which can be related to classical correlations and also to the intricate issues of ergodicity and mixing, an attempt was made to classify quantum dynamics similar to their classical counterpart. Of central interest in this study is the correlation function constructed in such a way that it has a well-defined classical limit. In this paper, we will study these quantities with their original formulation and without necessarily considering the classical limits to illuminate properties of certain quantum spin systems, such as the mixed-field Ising model and other models such as the Gaussian Orthogonal Ensemble (GOE) and Gaussian Unitary Ensemble (GUE) Rosenzweig Porter (RP) models. The mixed field Ising model is a simple model of $1/2-$ spins in a 1D chain with a nearest neighbour interaction and two transverse external fields. The model is known to exhibit both integrable and quantum chaotic limits Bañuls et al. (2011); Kim and Huse (2013a); Zhang et al. (2015); Roberts et al. (2015); Craps et al. (2020); Rodriguez-Nieva et al. (2023a); Camargo et al. (2024), depending on the strength of the transverse field. The eigenvalues of the model follow the RMT behaviour, yet the model does not have the expected chaotic dynamics, and exhibit weak chaos Fine et al. (2014); Craps et al. (2020); Kukuljan et al. (2017); Khemani et al. (2018); Xu and Swingle (2019). The RP model is a simple one-parameter model of random matrices that has a rich structure and is well studied in literature Rosenzweig and Porter (1960); Kravtsov et al. (2015); Venturelli et al. (2023); Zhang et al. (2023); De Tomasi et al. (2019); Buijsman (2024); Pino et al. (2019); Cadez et al. (2024); Bhattacharjee and Nandy (2024) for its properties. This simple one-parameter model remarkably has a fractal phase (an extended phase) apart from the usually found ergodic (chaotic) and localized (integrable) phases, depending on the value of its parameters. Our findings suggest that even if the eigenvalue statistics follow RMT behaviour it is still possible that the eigenvector correlations do not follow the same, as we will see in the case of the mixed-field Ising model. Such conclusions have also been made for another simple model of quantum chaos, Baker’s map Balazs and Voros (1989, ); O’Connor et al. (1992); Lakshminarayan (2019). Furthermore, for models like the RP model, we show that the average and variance of relaxation fluctuations can used to detect the transition quite well.
@@ II Setup and Random Matrix Theory
Thus, $c_{\text{avg}}$ is a measure of the distribution of eigenvectors in the two sub-spaces $A$ and $B$. Another feature of the average of correlations is that they capture the amount of spread of eigenstates in the subspace $A$ and $B$. If they are equally distributed then the average tends towards the classical time average and is equal to 1. This progression of results towards classically results means the result in the large $D$ limit (notice $\hbar_{\text{eff}}$ is defined to be equal to $\frac{1}{D}$ for such systems and large $D$ implies $\hbar_{\text{eff}}\rightarrow 0$). To capture this feature it is important to look at the scaling of $c_{\text{avg}}$ with $D$. In Lakshminarayan and Balazs (1994); Lakshminarayan (1998, 1997) it was suggested via various different models that $c_{\text{avg}}<1$, which implies the reluctance to participate equally in both partitions. However, there is still a possibility of $c_{\text{avg}}=1$ if there are symmetries present in the systems for e.g. the parity symmetry which forces the wavefunctions to be essentially identical in the two sub-spaces $A$ and $B$. Similarly the variance of $c(t)$ can be written as follows after simplification
@@
where $\braket{\cdot}$ denotes the ensemble average of the quantities. It is important to note that the above results are independent of the fraction $f$. We have plotted the above results in Fig. (1) along with the numerical results obtained using the GOE matrices. We have also plotted the results for the GUE matrices. For numerical purposes for both GOE and GUE type random matrices, we have taken $f=\frac{1}{2}$ and system sizes $D=128(2048),256(1024),512(512),1024(128),2048(64)$ ,4096(32),4096(32), 4096 ( 32 ), where the numbers in the bracket denote the number of samples. As outlined in Appendix A, we have the following results for GUE matrices
@@ III Baker’s map: A revisit
The quantized version of the map corresponds to the following unitary operator $U$, in position representation, with Hilbert space dimension $D$
@@
$(G_{D})_{l,m}=\frac{1}{\sqrt{D}}\exp\left(-2\pi i(l+\alpha)(m+\beta)\right)$
and different value of $\alpha$ and $\beta$ corresponds to same classical limit with different boundary condition Saraceno (1990); Lakshminarayan (2018). It was found in Saraceno (1990) that the anti-periodic boundary condition $\alpha=\beta=1/2$, chosen in this paper, maximally preserves the classical reflection symmetry. The classical map can be shown to be chaotic and mixing with the Lyapunov exponent equal to $\ln(2)$, which is also equal to the Lyapunov exponent obtained using OTOC in the quantized Baker’s map too Lakshminarayan (2019). We remark that we can form various kinds of generalized Baker’s map similar to Eq. (12), by taking into account the different partition schemes in the classical version, e.g. Eq. (12) is isomorphic to $(1/2,1/2)$ Bernoulli process Saraceno (1990). As another example, we can consider a classical map with Bernoulli scheme $(2/3,1/3)$ Saraceno (1990); Lakshminarayan (1997, 1998) given as follows
@@
The $(1/2,1/2)$ scheme also has a parity symmetry with the parity operator given by $P=-G_{D}^{2}$. The eigenvalues of the unitary operator $U$ are given by $e^{i\theta}$, where $\theta$ is the eigenangle, $\theta\in[0,2\pi]$ which are uniformly distributed on a circle. The spacing distribution of eigenangles for the chaotic operator follows the GOE distribution. In Fig. (2) we show the eigenangle spacing distribution of Baker’s map with Bernoulli scheme $(1/2,1/2)$ after symmetry resolution and unfolding. We observe quite a good agreement with the GOE spacing distribution, which is given by the red solid line in Fig. (2). We now study the $c_{avg}$ and $\sigma_{c}^{2}$ of the fluctuations for various schemes of Baker’s map.
@@
Figure 3: (Top) Logarithmic deviation of $c_{avg}$ from unity for quantum Baker’s map for various Bernoulli schemes. Notice that for the $(1/2,1/2)$ scheme, $c_{\text{avg}}=1$ and the corresponding data is not shown. We plot the logarithmic deviation of $c_{\text{avg}}$ from unity. (Bottom) The variance of fluctuation for quantum Baker’s amp for various Bernoulli schemes. We show the variation of the logarithmic variance with $\log(D)$. In both the plots the dashed lines are the best-fit lines obtained using sizes $D>200$.
The Fig (Bottom). (3) suggest the following trend of $c_{\text{avg}}$ and $\sigma_{c}^{2}$ with system size $D$
@@
For various schemes, we obtain the following values of the best-fit parameter for $c_{\text{avg}}=$
@@ IV Mixed-Field Ising model
for $k\in\{x,y,z\}$. We consider open boundary conditions, with $N$ sites with the Hilbert space of dimension $D=2^{N}$.
The model is known to exhibit both chaotic and integrable regimes Bañuls et al. (2011); Kim and Huse (2013a); Zhang et al. (2015); Roberts et al. (2015); Craps et al. (2020); Rodriguez-Nieva et al. (2023a); Camargo et al. (2024) depending on the strength of longitudinal and transverse fields as follows
@@
The above chaotic parameters are called the Banuls-Cirac-Hastings (BCH) parameters Bañuls et al. (2011). We would like to note that there are other parameters in the literature that are claimed to be maximally chaotic based on various other measures of chaos. Some of these parameters are Kim-Huse parameters Kim and Huse (2013b) and other parameters suggested in Rodriguez-Nieva et al. (2023b). We however choose the former, the BCH parameters and present the result for Kim-Huse parameters in Appendix B. Since the model has parity symmetry hence we look for these properties by focusing on symmetry resolves blocks, so as to eradicate the effect of symmetry on these quantities. Once the symmetries are resolved we then carry out our analysis.
@@
[…]
@@
with best-fit parameters $\kappa_{+}=0.46,\,\delta_{+}=1.54$. Unlike the behaviour $c_{\text{avg}}$ for the variance, we see that the scaling of variance with $D$. It has been observed in Lakshminarayan (1998, 1997) as well that for integrable systems $\sigma_{c}^{2}$ can scale with $D$ but the coefficient is non-universal with a strong dependence on partition $f$.
@@
where $\pm$ denotes the $+1$ and $-1$ parity sectors respectively with $\alpha_{\pm}\approx 1,\,\beta_{\pm}\approx\frac{7}{10}$, which deviates from the RMT predictions as discussed in Section II, where $\alpha=2$ and $\beta=1$. This implies that in some sense, the Ising model is less chaotic as compared to other models of quantum chaos which follow the RMT behaviour closely. We have already observed similar features in the case of the quantum Baker’s map. These results hint towards a hierarchy amongst systems exhibiting quantum chaos and suggest that not all systems are as chaotic as others. Furthermore, it shows that eigenvalue statistics are not enough to characterize the dynamics of the quantum systems, as it is possible that although the eigenvalue statistics of these systems follow RMT behaviour closely, the correlation of eigenvectors as discussed in the present context is not in agreement with the corresponding RMT results. We would also like to remark that spin $1/2$ systems are the most quantum systems we can have and thus they have peculiar features within the context of quantum chaos as has been reported before in various studies using other measures as well Fine et al. (2014); Craps et al. (2020); Kukuljan et al. (2017); Khemani et al. (2018); Xu and Swingle (2019).
@@ V Rosenzweig-Porter Model
Figure 7: The variation of $\braket{c(t)}$ with time $t$ corresponding to three different values of $\gamma$ in the RP model, Eq. (26). The three values correspond to three different regimes ergodic ( black), fractal (blue) and red (integrable). Results are shown for $D=512$ with $256$ disordered realization.
@@
We further study the properties of $\braket{c_{\text{avg}}}$ and $\braket{\sigma_{\text{c}}^{2}}$ for these three regimes with their proper scaling with size $D$, thus giving us the proper insight into the behaviour of these three regimes. In Fig. (8) (Top) we have numerically studied the variation of the average of the relaxation fluctuation with $\gamma$ for the GOE - RP model. We observe that for small $\gamma$ the model behaves like a GOE-like random matrix and transitions to a localized phase for large $\gamma$ via the fractal phase, where it eventually goes to zero. We have plotted these variations for various system sizes, $D=128(1024),256(512),512(256),1024(128),2048(64)$ and $4096(32)$ with the corresponding realizations shown in the brackets, in Fig. (8). For the $\braket{c_{\text{avg}}}$ we observe that there is a crossing at $\gamma=2$, capturing the transition from the ergodic to integrable regime. We further observe that the average relaxation fluctuation is incapable of detecting the fractal phase. However, if we study the variance of relaxation fluctuations it is seen there is a clear distinction between the three phases. In Fig. (8) (Bottom) we have plotted the variation of variation times the system size with $\gamma$. We observe that in both the ergodic (chaotic) as well the localized (integrable) regime the curves collapse. It is also observed that the $D^{2}\braket{\sigma^{2}_{c}}$ is constant till $\gamma\approx 1$, i.e. in the ergodic regime and then gradually increases till $\gamma\approx 2$ in the fractal phase and then further decreases and eventually reaches zero in the deep localized regime.
@@
For the ergodic ( chaotic) and localized (integrable) these exponents are in accordance with the analytical results from RMT. As we stated earlier in Sec. II, $c_{\text{avg}}$ captures the amount of spread of the wavefunctions in the two subspaces $A$ and $B$. For a system with chaos, it would mean that this quantity is as close to 1 as possible as $D$ becomes larger and larger. If it is less than 1 then it implies that spread is less and there is localization a fact well observed in previous studies as well Lakshminarayan and Balazs (1994). To be precise $\beta=0$ would imply a complete localization and the intermediate values $10<\beta<$ 1 would imply a fractal regime.
@@
As a further study, we also consider the GUE- RP model, which is given by Eq. (26) except now the matrix $B$ is a GUE matrix, whose elements, both real and imaginary parts, are drawn from a normal distribution with zero mean (for both diagonal and off-diagonal elements ) and variance is 1 and 1/2 for diagonal and off-diagonal elements respectively. It is also to be noted that for such a matrix the diagonal elements are real and off-diagonal elements are complex. The model exhibits a similar ergodic to localized transition with the fractal regime in between with corresponding transitions at $\gamma=1\,\text{and}\,2$ respectively.
@@
Finally, we revert back to the variation of $\braket{c_{\text{avg}}}$ with $D$. Since we have an intersection of curves at $\gamma=2$, we can perform a data collapse. Similar to the $\braket{r}$ value Buijsman (2024); Pino et al. (2019), we find that for $\braket{c_{\text{avg}}}$ too has the scaling collapse that can be obtained by plotting it against $(\gamma-2)\ln(D)$ for both GOE and GUE-RP model. We show this in Fig. 11. All these results exploring the correlations of eigenvectors in this system give an alternative way to study the chaotic-integrable transition along with an alternative and novel way to explore the properties of the fractal states as well. The possibility to use $\braket{c_{\text{avg}}}$ and $\sigma^{2}$ to identify such transition was already pointed out in Lakshminarayan and Balazs (1994), based on the study of Standard map.

### v3

@@ I Introduction
With the motivation to study quantum chaotic systems and possibly to classify them, similar to their classical counterpart, we utilize the ideas put forth in Lakshminarayan and Balazs (1994); Lakshminarayan (1997, 1998) regarding the relaxation fluctuations of the correlation function and its intrinsic connection to chaos. The study was motivated by and studied using models having well-defined classical limits which are known to be chaotic such as in quantum maps Haake (1991); Lakshminarayan (2018); Balazs and Voros (1989); Lakshminarayan and Balazs (1994); Lakshminarayan (1997, 1998). By studying the quantum correlation functions which can be related to classical correlations and also to the intricate issues of ergodicity and mixing, an attempt was made to classify quantum dynamics similar to their classical counterpart. Of central interest in this study is the correlation function constructed in such a way that it has a well-defined classical limit. In this paper, we will study these quantities with their original formulation and without necessarily considering the classical limits to illuminate properties of certain quantum spin systems, such as the mixed-field Ising model and other models such as the Gaussian Orthogonal Ensemble (GOE) and Gaussian Unitary Ensemble (GUE) Rosenzweig Porter (RP) models. The mixed field Ising model is a simple model of $1/2-$ spins in a 1D chain with a nearest neighbour interaction and two transverse external fields. The model is known to exhibit both integrable and quantum chaotic limits Bañuls et al. (2011); Kim and Huse (2013); Zhang et al. (2015); Roberts et al. (2015); Craps et al. (2020); Rodriguez-Nieva et al. (2023); Camargo et al. (2024), depending on the strength of the transverse field. The eigenvalues of the model follow the RMT behaviour, yet the model does not have the expected chaotic dynamics, and exhibit weak chaos Fine et al. (2014); Craps et al. (2020); Kukuljan et al. (2017); Khemani et al. (2018); Xu and Swingle (2019). The RP model is a simple one-parameter model of random matrices that has a rich structure and is well studied in literature Rosenzweig and Porter (1960); Kravtsov et al. (2015); Venturelli et al. (2023); Zhang et al. (2023); De Tomasi et al. (2019); Buijsman (2024); Pino et al. (2019); Cadez et al. (2024); Bhattacharjee and Nandy (2024) for its properties. This simple one-parameter model remarkably has a fractal phase (an extended phase) apart from the usually found ergodic (chaotic) and localized (integrable) phases, depending on the value of its parameters. Our findings suggest that even if the eigenvalue statistics follow RMT behaviour it is still possible that the eigenvector correlations do not follow the same, as we will see in the case of the mixed-field Ising model. Such conclusions have also been made for another simple model of quantum chaos, Baker’s map Balazs and Voros (1989, ); O’Connor et al. (1992); Lakshminarayan (2019). Furthermore, for models like the RP model, we show that the average and variance of relaxation fluctuations can used to detect the transition quite well.
@@ II Setup and Random Matrix Theory
Thus, $c_{\text{avg}}$ is a measure of the distribution of eigenvectors in the two sub-spaces $A$ and $B$. Another feature of the average of correlations is that they capture the amount of spread of eigenstates in the subspace $A$ and $B$. If they are equally distributed then the average tends towards the classical time average and is equal to 1. This progression of results towards classically results means the result in the large $D$ limit (notice $\hbar_{\text{eff}}$ is defined to be equal to $\frac{1}{D}$ for such systems and large $D$ implies $\hbar_{\text{eff}}\rightarrow 0$). To capture this feature it is important to look at the scaling of $c_{\text{avg}}$ with $D$. In Lakshminarayan and Balazs (1994); Lakshminarayan (1998, 1997) it was suggested via various different models that $c_{\text{avg}}<1$, which implies the reluctance to participate equally in both partitions. However, there is still a possibility of $c_{\text{avg}}=1$ if there are symmetries present in the systems. One such case where this can happen is, the Baker’s map in the presence of parity symmetry. The presence of parity symmetry forces the wave function to be essentially identical in the two subspaces thus giving $c_{avg}=1$ Lakshminarayan (1997). For this reason for all the studies in the paper, we remove all the unitary symmetries ( for e.g. parity ) associated with the unitary evolution operator $U$ or Hamiltonian $H$ and always work out the details of $c_{avg}$ and $\sigma_{c}$ using the symmetry resolved blocks. Similarly the variance of $c(t)$ can be written as follows after simplification
@@
where $\braket{\cdot}$ denotes the ensemble average of the quantities. It is important to note that the above results are independent of the fraction $f$. We have plotted the above results in Fig. (1) along with the numerical results obtained using the GOE matrices. We have also plotted the results for the GUE matrices. For numerical purposes for both GOE and GUE type random matrices, we have taken $f=\frac{1}{2}$ and system sizes $D=128(2048),256(1024),512(512),1024(128),2048(64)$ ,4096(32),4096(32), 4096 ( 32 ), where the numbers in the bracket denote the number of realizations taken for averaging. As outlined in Appendix A, we have the following results for GUE matrices
@@ III Baker’s map: A revisit
The quantized version of the map corresponds to the following unitary operator $U$ Balazs and Voros (1989), in position representation, with Hilbert space dimension $D$
@@
$(G_{D})_{l,m}=\frac{1}{\sqrt{D}}\exp\left(-2\pi i(l+\alpha)(m+\beta)\right)$ (14)
and different value of $\alpha$ and $\beta$ corresponds to same classical limit with different boundary condition Saraceno (1990); Lakshminarayan (2018). It was found in Saraceno (1990) that the anti-periodic boundary condition $\alpha=\beta=1/2$, chosen in this paper, maximally preserves the classical reflection symmetry. The classical map can be shown to be chaotic and mixing with the Lyapunov exponent equal to $\ln(2)$, which is also equal to the Lyapunov exponent obtained using OTOC in the quantized Baker’s map Lakshminarayan (2019). We remark that we can construct various other kinds of generalized Baker’s map similar to Eq. (12), by taking into account the different partition schemes in the classical version, e.g. Eq. (12) is isomorphic to $(1/2,1/2)$ Bernoulli process Saraceno (1990). As another example, we can consider a classical map with Bernoulli scheme $(2/3,1/3)$ Saraceno (1990); Lakshminarayan (1997, 1998) given as follows
@@
The $(1/2,1/2)$ scheme also has a parity symmetry with the parity operator given by $P=-G_{D}^{2}$. The eigenvalues of the unitary operator $U$ are given by $e^{i\theta}$, where $\theta$ is the eigenangle, $\theta\in[0,2\pi]$ which are uniformly distributed on a circle. The spacing distribution of eigenangles for the chaotic operator follows the GOE distribution. In Fig. (2) we show the eigenangle spacing distribution of Baker’s map with Bernoulli scheme $(1/2,1/2)$ after symmetry resolution and unfolding. We observe quite a good agreement with the GOE spacing distribution, which is given by the red solid line in Fig. (2). We now study the $c_{avg}$ and $\sigma_{c}^{2}$ of the fluctuations for various schemes of Baker’s map with $f=1/2$. As an important side note we would like to mention that another way to work without parity symmetry in Baker’s map is to choose $\alpha\,\text{or}\,\beta\neq 1/2$ in Eq. (14) thus explicitly breaking the parity symmtery.
@@
Figure 3: (Top) Logarithmic deviation of $c_{avg}$ from unity for quantum Baker’s map for various Bernoulli schemes. Notice that for the $(1/2,1/2)$ scheme, $c_{\text{avg}}=1$ and the corresponding data is not shown. We plot the logarithmic deviation of $c_{\text{avg}}$ from unity. (Bottom) The variance of fluctuation for quantum Baker’s amp for various Bernoulli schemes. We show the variation of the logarithmic variance with $\ln(D)$. In both the plots the dashed lines are the best-fit lines obtained using sizes $D>200$. The green line is the analytical result given by Eq. (9).
The Fig. (3) (Bottom) suggest the following trend of $c_{\text{avg}}$ and $\sigma_{c}^{2}$ with system size $D$
@@
For various schemes, we obtain the following values of the best-fit parameter for $c_{\text{avg}}$
@@ IV Mixed-Field Ising model
for $k\in\{x,y,z\}$. We consider open boundary conditions, with $N$ sites with the Hilbert space of dimension $D=2^{N}$. For all the numerical computations we choose $f=1/2$.
The model is known to exhibit both chaotic and integrable regimes Bañuls et al. (2011); Kim and Huse (2013); Zhang et al. (2015); Roberts et al. (2015); Craps et al. (2020); Rodriguez-Nieva et al. (2023); Camargo et al. (2024) depending on the strength of longitudinal and transverse fields as follows
@@
The above chaotic parameters are called the Banuls-Cirac-Hastings (BCH) parameters Bañuls et al. (2011). We would like to note that there are other parameters in the literature that are claimed to be maximally chaotic based on various other measures of chaos. Some of these parameters are Kim-Huse parameters Kim and Huse (2013) and other parameters suggested in Rodriguez-Nieva et al. (2023). We however choose the former, the BCH parameters and present the result for Kim-Huse parameters in Appendix B. Since the model has parity symmetry hence we look for these properties by focusing on symmetry resolves blocks, so as to eradicate the effect of symmetry on these quantities. Once the symmetries are resolved we then carry out our analysis. We also note that for the present case, the projection operators are diagonal in the $\sigma_{z}$ and has been chosen accordingly. This indicates, that all the results, including the RMT predictions in section II, have a basis dependence, in the sense that we choose a basis such that the projection operators are diagonal.
@@
The integrable and chaotic regimes are quite evident from the scaling of $(c_{\text{avg}})_{+}$ and $\sigma_{+}^{2}$ with $D$, where + sign denotes the +1 parity sector. For the integrable case, the results in Fig. (5), suggest the scaling behaviour of the following form
@@
with best-fit parameters $\kappa_{+}=0.46,\,\delta_{+}=1.54$. Unlike the behaviour $c_{\text{avg}}$ for the variance, we see that the scaling of variance with $D$. It has been observed in Lakshminarayan (1998, 1997) as well that, for integrable systems $\sigma_{c}^{2}$ can scale with $D$ but the coefficient is non-universal with a strong dependence on partition $f$.
@@
where $\pm$ denotes the $+1$ and $-1$ parity sectors respectively with $\alpha_{\pm}\approx 1,\,\beta_{\pm}\approx\frac{7}{10}$, which deviates from the RMT predictions as discussed in Section II, where $\alpha=2$ and $\beta=1$. This implies that in some sense, the Ising model is less chaotic as compared to other models of quantum chaos which follow the RMT behaviour closely. We have already observed similar features in the case of the quantum Baker’s map. These results hint towards a hierarchy among systems exhibiting quantum chaos and suggest that not all systems are as chaotic as others. Furthermore, it shows that eigenvalue statistics are not enough to characterize the dynamics of the quantum systems, as it is possible that although the eigenvalue statistics of these systems follow RMT behaviour closely, the correlation of eigenvectors as discussed in the present context is not in agreement with the corresponding RMT results. We would also like to remark that spin $1/2$ systems are the most quantum systems we can have and thus they have peculiar features within the context of quantum chaos as has been reported before in various studies using other measures as well Fine et al. (2014); Craps et al. (2020); Kukuljan et al. (2017); Khemani et al. (2018); Xu and Swingle (2019).
@@ V Rosenzweig-Porter Model
Figure 7: The variation of $\braket{c(t)}$ with time $t$ corresponding to three different values of $\gamma$ in the RP model, Eq. (27). The three values correspond to three different regimes ergodic ( black), fractal (blue) and red (integrable). Results are shown for $D=512$ with $256$ disordered realizations.
@@
We further study the properties of $\braket{c_{\text{avg}}}$ and $\braket{\sigma_{\text{c}}^{2}}$ for these three regimes with their proper scaling with size $D$ and $f=1/2$, thus giving us the proper insight into the behaviour of these three regimes. In Fig. (8) (Top) we have numerically studied the variation of the average of the relaxation fluctuation with $\gamma$ for the GOE - RP model. We observe that for small $\gamma$ the model behaves like a GOE-like random matrix and transitions to a localized phase for large $\gamma$ via the fractal phase, where it eventually goes to zero. We have plotted these variations for various system sizes, $D=128(1024),256(512),512(256),1024(128),2048(64)$ and $4096(32)$ with the corresponding realizations shown in the brackets, in Fig. (8). For the $\braket{c_{\text{avg}}}$ we observe that there is a crossing at $\gamma=2$, capturing the transition from the ergodic to integrable regime. We further observe that the average relaxation fluctuation is incapable of detecting the fractal phase. However, if we study the variance of relaxation fluctuations it is seen there is a clear distinction between the three phases. In Fig. (8) (Bottom) we have plotted the variation of variation times the system size with $\gamma$. We observe that in both the ergodic (chaotic) as well the localized (integrable) regime the curves collapse. It is also observed that the $D^{2}\braket{\sigma^{2}_{c}}$ is constant till $\gamma\approx 1$, i.e. in the ergodic regime and then gradually increases till $\gamma\approx 2$ in the fractal phase and then further decreases and eventually reaches zero in the deep localized regime.
@@
For the ergodic ( chaotic) and localized (integrable) these exponents are in accordance with the analytical results from RMT. As we stated earlier in Sec. II, $c_{\text{avg}}$ captured the amount of spread of the wavefunctions in the two subspaces $A$ and $B$. For a system with chaos, it would mean that this quantity is as close to 1 as possible as $D$ becomes larger and larger. If it is less than 1 then it implies that spread is less and there is localization a fact well observed in previous studies as well Lakshminarayan and Balazs (1994). To be precise $\beta=0$ would imply a complete localization and the intermediate values $10<\beta<$ 1 would imply a fractal regime.
@@
As a further study, we also consider the GUE- RP model, which is given by Eq. (27) except now the matrix $B$ is a GUE matrix, whose elements, both real and imaginary parts, are drawn from a normal distribution with zero mean (for both diagonal and off-diagonal elements ) and variance is 1 and 1/2 for diagonal and off-diagonal elements respectively. It is also to be noted that for such a matrix the diagonal elements are real and off-diagonal elements are complex. The model exhibits a similar ergodic to localized transition with the fractal regime in between with corresponding transitions at $\gamma=1$and  2 respectively.
@@
Finally, we revert back to the variation of $\braket{c_{\text{avg}}}$ with $D$. Since we have an intersection of curves at $\gamma=2$, we can perform a data collapse. Similar to the $\braket{r}$ value Buijsman (2024); Pino et al. (2019), we find that for $\braket{c_{\text{avg}}}$ too has the scaling collapse that can be obtained by plotting it against $(\gamma-2)\ln(D)$ for both GOE and GUE-RP model. We show this in Fig. 11. All these results exploring the correlations of eigenvectors in this system give an alternative way to study the chaotic-integrable transition along with an alternative way to explore the properties of the fractal states as well. The possibility to use $\braket{c_{\text{avg}}}$ and $\sigma^{2}$ to identify such transition was already pointed out in Lakshminarayan and Balazs (1994), based on the study of Standard map.